OpenStudy (anonymous):

What is the equation of a line that passes through the point (0, -3) and is perpendicular to a line with a slope of 4

OpenStudy (triciaal):

for perpendicular lines the product of the slopes is -1 let the slopes = m1 and m2 then m1*m2 =-1 so you have a slope of 4 what is the the slope of the perpendicular line?

OpenStudy (triciaal):

next you have the slope of the line, m and 2 points on the line (x, y) and (0, -3) the equation of the line is (y-y1) = m(x- x1) where (x,y) is any point on the line and (x1, y1) is a known point on the line

OpenStudy (triciaal):

any questions?